OpenStudy (anonymous):

Identify which equation is a quadratic equation. y = −6x − 3 y + x2 = (x − 1)(x + 7) y − 9 = (x − 1)(x2 + 4) y = 5x(x + 3) − 8

OpenStudy (anonymous):

Idk how to determine the difference between a regular and quadratic equation

OpenStudy (dumbcow):

quadratic just means the highest exponent is 2 so if the equation has an "x^2" term , then it is a quadratic

OpenStudy (dumbcow):

looks like they are trying to trick you here....make sure you combine all like terms, try to get y by itself on left side

OpenStudy (anonymous):

sooo y + x2 = (x − 1)(x + 7) y − 9 = (x − 1)(x2 + 4) coul be the answers?

OpenStudy (anonymous):

i think its this one : y + x2 = (x − 1)(x + 7)

OpenStudy (dumbcow):

first one could be but what happens after you multiply (distribute) and combine like terms? second one is not because what happens when you multiply x by x^2